I have a problem that I can't seem to be able to set up. It is Pablo is 3 miles behind Marc. Pablo increases his speed to 5 miles per hour, while Marc, runs at 4mph. At this rate, how many miles will Pablo have to run to catch up to Marc? How long will it take?
Answer by ankor@dixie-net.com(22740)

Pablo is 3 miles behind Marc.
Pablo increases his speed to 5 miles per hour, while Marc, runs at 4mph.
At this rate, how many miles will Pablo have to run to catch up to Marc?
How long will it take?
:
Let t = time for P to catch M
P has to run 3 more miles than M
Write a distance equation; dist = speed * time
5t = 4t + 3
5t - 4t = 3
t = 3 hrs for P to catch M
:
How many miles will P have to run to catch M
3 * 5 = 15 miles
:
:
Check this, find the distance each runs
5*3 = 15 mi
4*3 = 12 mi
------------
diff: 3 mi